Calculate the unemployment rate when the number of unemployed is 1.5 million and the total labor force is 34 million.


1
Expert's answer
2021-02-02T14:47:00-0500

Unemployment Rate can be calculated as follows:


"U=\dfrac{Unemployed\ People}{Labor\ Fprce}\times100\%,""U=\dfrac{1500000}{34000000}\times 100\%=4.41\%."